Do you know compounding interest; Pe rt; find the "r" rate of interest
At what interest rate must 4800 be compounded annually to equal 9051.12 after 13 years, trying to find a formula like this in one of my 6 books, and I could find P, T, A but not "r",
raysecrets @hotmail.com, I have been working on this homework problem for 2 days now. Any help will be great.

Answer by scott8148(6628)

FV=P(1+r)^n
__ FV is the future value
__ P is the starting amount (principal)
__ r is the interest rate per compounding period (as a decimal)
__ n is the number of compounding periods

the "Pe rt" is for continuous compounding __ e is the base for natural logs ( Euler's number)

9051.12=4800(1+r)^13 __ dividing by 4800 __ 9051.12/4800=(1+r)^13

taking the log __ log(9051.12/4800)=13(log(1+r))

dividing by 13 __ (log(9051.12/4800))/13=log(1+r)

taking antilog __ 10^((log(9051.12/4800))/13)=1+r

subtracting 1 __ (10^((log(9051.12/4800))/13))-1=r